Mon 30 Nov - Eduroam WiFi Connectivity To WiNet, and Telephone Problems
Eduroam Connectivity To WiNet
Users can connect to the Wireless Network using Eduroam but may experience problems authenticating to some network resources.
Telephone Problems: Dialing numbers beginning 9 from extensions beginning with a 1
There is currently an intermittent fault affecting extensions beginning with a 1. When you try to dial a number beginning with a 9 the system returns a "Your call cannot be completed as dialed" error message. ISS are investigating the problem and apologise for any inconvenience. In order to reach other extensions beginning with a 9 callers can dial the full DDI number beginning 8015245(9xxxx), where 9xxxx is the 5 digit extension number you wish to dial.
Fri 27 Nov - All systems ran without problems
Thu 26 Nov - Eduroam Service Unavailable & HPC Unavailable Due to Brownout
Eduroam Service Unavailable
Eduroam continues to be unavailable. Other services requiring authentication are working. ISS will restore the Eduroam service as soon as possible.
HPC Unavailable Due to Brownout
Due to a brownout overnight, power was lost to two racks in the HPC cluster. The affected nodes were rebooted and are usable again.
Wed 25 Nov - Problem With System Authentication, & Room Booking and Academic Timetable: Enterprise System Upgrade
Problem With System Authentication
VPN, Eduroam and other systems that require authentication were unavailable at 'login'. If you are already logged in you will be able to continue using the system. UPDATE at19:27: The VPN service is now available again, ISS will continue to monitor the service closely.
Room Booking and Academic Timetable: Enterprise System Upgrade
The Enterprise Room Booking system is being upgraded on the 24th – 25th November.
Tue 24 Nov - Room Booking and Academic Timetable: Enterprise System Upgrade
The Enterprise Room Booking system is being upgraded on the 24th – 25th November. Please be aware that the following timetable services will be unavailable during this time: Web Room Booking, Room Timetable. Please make sure all bookings for the above dates have been made by the 23rd November. Academic Timetable information will still be displayed in LUSI and students will still be able to view their timetable through LUVLE. We apologise for any inconvenience this may cause.
Mon 23 Nov - H: Drive Problem, and Fake Adobe PDF Reader Message
H: Drive Problem
Some users were experiencing problems saving files on H: Drive. ISS located the issue and resolved it.
Fake Adobe PDF Reader Message
A fake email, subject: "Update New Adobe PDF Reader For Windows" has been seen. Do NOT follow the link and do NOT download the software.
Sat 07 to Sat 21 Nov - ISS Staff moved to the new ISS Building.
To minimise the disruption to services, the move was phased over three weekends.
The new ISS building has access control, as befits a university data centre. If you need to visit staff at the new building, please use the telephone in the Building’s lobby and call the member of ISS staff that you are visiting – they can then meet you at the inner door.
The ISS Service Desk (for those visiting in person) has already moved to the Learning Zone, so will be unaffected by these staff moves. Service Desk staff responding to e-mail and telephone enquiries will be involved in the move, but will attempt to maintain levels of service where possible. Where possible, please would you postpone non-urgent enquiries from Friday afternoon and Monday morning to permit staff to complete the move to their new locations.
Please note that the new telephone numbers for Learning Technology (they range from 10651 to 10657) are as published in this year’s University Telephone Directory; Learning Technology’s former telephone numbers will continue to work for the next 6 months.
Fri 20 Nov - Webmail Upgrade Phish Attack
Another fake message, subject: "Webmail Upgrade" is circulating. Do NOT follow the link in the message. Do NOT reply to the message
Thu 19 Nov - All systems ran without problems
Wed 18 Nov - All systems ran without problems
Tue 17 Nov - Network at Risk - Engineering and Management School, and Fake Email Message "Your mailbox has been deactivated"
Network at Risk - Engineering and Management School
07:00-08:30: During the 'at risk' period on Tuesday 17th November two switches were swapped in Engineering and the Management school. Work took no more than 10 minutes and affected a maximum of 50 computers throughout the work.
Fake Email Message "Your mailbox has been deactivated"
Fake messages subject: "your mailbox has been deactivated" are arriving on-campus. The message has not been sent from ISS. Do not open the attachment. If you have opened the attachment contact the ISS Service Desk immediately.
Mon 16 Nov - All systems ran without problems
Sat 14 Nov - ISS Staff Move to ISS Building (move 2 of 3)
The second of 3 weekend staff moves to the ISS Building were undertaken:
- Sat 07 Nov CIS Academic (part), CIS LTG (part), TIG Systems, TIG Networking Stores
Fri 13 Nov - All systems ran without problems
Thu 12 Nov - All systems ran without problems
Wed 11 Nov - Problems with AV Equipment in Frankland Lecture Theatre
There were a few problems in the Frankland lecture theatre with AV equipment where only video from the PC could be projected. Users wishing to play DVDs, use laptops or other equipment were required to make alternative arrangements. ISS engineers worked tirelessly to rectify the fault with the minimum of delay.
Tue 10 Nov - All systems ran without problems
Mon 09 Nov - All systems ran without problems
Sat 07 Nov - ISS Staff Move to ISS Building (move 1 of 3)
The first of 3 weekend staff moves to the ISS Building were undertaken:
- Sat 07 Nov ISS Director John Gallagher, Departmental Office, CIS Resources, CIS Database Administration, TIG Networking (part), and USG (from the Library)
Fri 06 Nov - All systems ran without problems
Thu 05 Nov - All systems ran without problems
Wed 04 Nov - LUVLE Maintenance
15:00-16:00: Parts of LUVLE were inaccessible on Wed 4th November due to maintenance work. LUVLE spaces affected include those belonging to Accounting & Finance, Biology, CEEC, CELT, Health Research, Economics, Entrepreneurship, Environmental Science, History, Marketing, Management, Management Science, Organisation Work and Technology, Physics and Psychology.
Tue 03 Nov - All systems ran without problems
Mon 02 Nov - HPC service restored
H: Drive is now fully functional. At a future date maintenance will be scheduled to finalize the solution to the issue. During this scheduled maintenance period users in area 18 will experience a short outage.
Service to the HPC has been restored after campus-wide power brown-outs over the weekend.